# Constrained Discrete Diffusion

### v2

@@ C.1 Toxicity Mitigation
To quantify the fluency and coherence of generated continuations we follow the increasingly adopted LLM-as-a-Judge evaluation. This method has been shown to result in similar evaluation as human judgment [64]. Specifically, each post-prefix sentence is scored by the instruction-tuned GEMMA-3-27B-IT model and issued the following prompt:
@@
As a baseline for toxicity reduction in natural language generation we include the PPLM methodology [26] . PPLM uses a frozen pretrained GPT-2 generator for output steering. At every decoding step the algorithm runs the usual forward pass, takes a few gradient-ascent steps on the cached key/value activations to maximize the score of a small attribute model which in this case is a toxicity classifier trained on the Jigsaw toxicity dataset. It simultaneously keeps the perturbed logits close to the original distribution with a KL penalty and a geometric-mean fusion term. We use the authors implementation and code structure without architectural modifications. The underlying steering mechanism, loss terms and hyper-parameter values are unchanged from the authors’ baseline.
@@
For another toxicity control baseline we additionally implemented FUDGE [15] on the Jigsaw Toxicity dataset. Following the authors implementation we use an LSTM followed by a linear output layer trained to predict future toxicity from partial prefixes. At inference time we add its scaled log non-toxicity probability to GPT-2’s next-token logits for the top candidates then sample from the re-weighted distribution. While the original implementation of FUDGE fixes the guidance strength $\lambda=1$, we perform a hyperparameter seach and sample from the resulting distribution with $\lambda\in\{1,\dots,9\}$, displaying the toxicity and perplexity results in Table 2. We use the authors’ original code structure, and base language model.
@@
While recent literature [8] has proposed adaptations of classifier-based guidance [28] and classifier-free guidance [65] for discrete diffusion models, these approaches have only been employed for settings with limited vocabulary sizes (e.g., molecule generation where the vocabulary is composed of 40 tokens). For natural language tasks which have much larger vocabulary sizes, for instance, the GPT2 tokenizer is composed of over 50,000 tokens–these techniques have not been applied. Currently, this is a scalability issue, as these approaches require that guidance vectors are computed of size $\mathbb{R}^{L\times N}$, making it much simpler and efficient when operating on a limited vocabulary size. Hence, as there is no prior literature to the extent of our knowledge that has scaled these techniques to natural language tasks, we defer comparison to these methods for the molecular generation setting.
@@
For the ALM projection, we initialize using the following hyperparameters: $\lambda_{\text{init}}=0.0,\mu_{\text{init}}=1.0,\mathrm{outer\_iter}_{\text{max}}=1000,\mathrm{inner\_iter}_{\text{max}}=10,\eta=0.20$. We choose these setting for the toxicity experiment as they result in the lowest perplexity in the Appendix D ablation study. For natural language toxicity reduction details we use a generation length of 100 tokens.
C.2 Molecular Generation
@@
For our molecule generation experiments, we utilize the QM9 dataset [51], as used by Schiff et al. [8]. This dataset comprises approximately 133,885 small organic molecules, each encoded as a SMILES string [48], which compactly represents the molecular structure through a sequence of characters. The SMILES representation facilitates discrete modeling by enabling the treatment of molecular generation as a sequence prediction task, making it particularly amenable to discrete diffusion approaches. By leveraging QM9, we can rigorously evaluate the performance of our generative models on tasks that require both the preservation of chemical validity and the precise control of molecular properties, aligning with established protocols in the literature.
@@
C.2.1 Synthetic Accessibility Constraints
@@
We train our surrogate model on the QM9 dataset [51] but manually label the training data with with RDKIT’s synthetic accessibility score [52]. We finetune GPT2 (124M) to act as this surrogate model and directly output a score $s\in[0,10]$.
@@
For this setting, we follow the FUDGE implementation provided by [8]. However, while their guidance is trained on QED scores, labeling samples from the QM9 dataset with these scores, we adapt it to label the training data with RDKIT’s computation of the synthetic accessibility scores.
@@
Similar to the implementation of FUDGE, we adapt the guidance mechanisms provided by [8] for QED by retraining on synthetic accessibility labels. Otherwise, our implementation mirrors their approach for QED guidance.
@@ C.2.2 Novelty Constraints
For this application, we similarly follow the FUDGE implementation and configuration in [8] for QED guidance.
@@
We use classifier based guidance and classifier free guidance for QED as implemented in [8].
C.3 Instruction Following
@@
To quantify the fluency and coherence of generated continuations we follow the increasingly adopted LLM-as-a-Judge evaluation. This method has been shown to result in similar evaluation as human judgment [64]. Specifically, each post-prefix sentence is scored by the instruction-tuned Gemini-2.0-Flash model and issued the following prompt:
@@
While in Section 5.1, we directly utilize the pretrained MDLM checkpoint provided by [7], for the experiments in Section 5.3, we pretrain and subsequently instruction-tune a slightly larger MDLM model, adopting the Llama 1 tokenizer. Empirical observations indicated that employing a marginally larger model and vocabulary size was essential for optimal performance on instruction-guided tasks. Consequently, we adopt this enhanced model for all subsequent evaluations in this context, including both the base MDLM experiments and the proposed CDD framework. Our instruction tuning is conducted using the ShareGPT dataset, and all pretraining follows the design choices and hyperparameters provided by [7].
C.3.1 Counting Constraints
@@ C.3.2 Lexical Constraints
We finetune models on a dataset of examples constructed from evaluation of GPT-4 on the COLLIE dataset provided by Yao et al. [53]. As GPT-4 performs well on these examples, we apply knowledge distillation by compiling these generations into an instruction tuning dataset. The dataset can be directly constructed from the logs included in the code base from Yao et al.
@@
Appendix D Hyper-parameter Ablation Study
@@
We include a constraint-start ablation, where we examine the runtime/perplexity tradeoff by turning the projection operator on only after specific amounts of reverse-diffusion steps. Early enforcements allows the projection to satisfy the constraints at very early iterations at the cost of increased runtime. Notably, across all setting, constraint satisfaction is met for all samples.
ALM hyper-parameters Perplexity Wall-clock time (s) LLM-Judge
@@
Table 3: Perplexity and wall-clock time statistics, together with LLM-Judge coherence scores, across all ALM hyper-parameter settings.
Appendix E Runtime Analysis
To benchmark the computational overhead of CDD as compared to the baselines, we compare the runtime of our method and the other baselines on the toxicity mitigation task. To normalize for the overhead of base models (MDLM and GPT2), we make two adjustment: we reduce the number of diffusion steps from 1000 to 100 and adjust the generation sequence length to 1024; prior work on discrete diffusion has indicated that these hyperparameters can be used to approximately match the inference time of autoregressive models [9]. Empirically, we find that, while the speed is not identical, these adjustments result in a much more similar tokens-per-second rate for the base models. Practically, MDLM could be accelerated even further by reducing the number of diffusion steps further, but we abstain from this so as to maintain consistency with the baselines. For every system we report the mean generation time, the 25th/50th/75th percentiles, and the slowdown relative to the respective baseline.
@@ Appendix E Runtime Analysis
[…]
@@ Appendix F Missing Proofs
[…]
@@
[…]
@@
[…]

### v3

@@ B.1 Toxicity Mitigation
To quantify the fluency and coherence of generated continuations we follow the increasingly adopted LLM-as-a-Judge evaluation. This method has been shown to result in similar evaluation as human judgment [zheng2023judgingllmasajudgemtbenchchatbot]. Specifically, each post-prefix sentence is scored by the instruction-tuned GEMMA-3-27B-IT model and issued the following prompt:
@@
As a baseline for toxicity reduction in natural language generation we include the PPLM methodology [dathathri2019plug] . PPLM uses a frozen pretrained GPT-2 generator for output steering. At every decoding step the algorithm runs the usual forward pass, takes a few gradient-ascent steps on the cached key/value activations to maximize the score of a small attribute model which in this case is a toxicity classifier trained on the Jigsaw toxicity dataset. It simultaneously keeps the perturbed logits close to the original distribution with a KL penalty and a geometric-mean fusion term. We use the authors implementation and code structure without architectural modifications. The underlying steering mechanism, loss terms and hyper-parameter values are unchanged from the authors’ baseline.
@@
For another toxicity control baseline we additionally implemented FUDGE [Yang_2021] on the Jigsaw Toxicity dataset. Following the authors implementation we use an LSTM followed by a linear output layer trained to predict future toxicity from partial prefixes. At inference time we add its scaled log non-toxicity probability to GPT-2’s next-token logits for the top candidates then sample from the re-weighted distribution. While the original implementation of FUDGE fixes the guidance strength $\lambda=1$, we perform a hyperparameter seach and sample from the resulting distribution with $\lambda\in\{1,\dots,9\}$, displaying the toxicity and perplexity results in Table 2. We use the authors’ original code structure, and base language model.
@@
While recent literature [schiff2024simple] has proposed adaptations of classifier-based guidance [ho2020denoising] and classifier-free guidance [ho2022classifier] for discrete diffusion models, these approaches have only been employed for settings with limited vocabulary sizes (e.g., molecule generation where the vocabulary is composed of 40 tokens). For natural language tasks which have much larger vocabulary sizes, for instance, the GPT2 tokenizer is composed of over 50,000 tokens–these techniques have not been applied. Currently, this is a scalability issue, as these approaches require that guidance vectors are computed of size $\mathbb{R}^{L\times N}$, making it much simpler and efficient when operating on a limited vocabulary size. Hence, as there is no prior literature to the extent of our knowledge that has scaled these techniques to natural language tasks, we defer comparison to these methods for the molecular generation setting.
@@
For the ALM projection, we initialize using the following hyperparameters: $\lambda_{\text{init}}=0.0,\mu_{\text{init}}=1.0,\mathrm{outer\_iter}_{\text{max}}=1000,\mathrm{inner\_iter}_{\text{max}}=10,\eta=0.20$. We choose these setting for the toxicity experiment as they result in the lowest perplexity in the Appendix C ablation study. For natural language toxicity reduction details we use a generation length of 100 tokens.
B.2 Molecular Generation
@@
For our molecule generation experiments, we utilize the QM9 dataset [ruddigkeit2012enumeration], as used by schiff2024simple. This dataset comprises approximately 133,885 small organic molecules, each encoded as a SMILES string [weininger1988smiles], which compactly represents the molecular structure through a sequence of characters. The SMILES representation facilitates discrete modeling by enabling the treatment of molecular generation as a sequence prediction task, making it particularly amenable to discrete diffusion approaches. By leveraging QM9, we can rigorously evaluate the performance of our generative models on tasks that require both the preservation of chemical validity and the precise control of molecular properties, aligning with established protocols in the literature.
@@
Figure 6: SA projection illustration.
B.2.1 Synthetic Accessibility Constraints
@@
We train our surrogate model on the QM9 dataset [ruddigkeit2012enumeration] but manually label the training data with RDKIT’s synthetic accessibility score [bento2020open]. We finetune GPT2 (124M) to act as this surrogate model and directly output a score $s\in[0,10]$.
@@
For this setting, we follow the FUDGE implementation provided by [schiff2024simple]. However, while their guidance is trained on QED scores, labeling samples from the QM9 dataset with these scores, we adapt it to label the training data with RDKIT’s computation of the synthetic accessibility scores.
@@
Similar to the implementation of FUDGE, we adapt the guidance mechanisms provided by [schiff2024simple] for QED by retraining on synthetic accessibility labels. Otherwise, our implementation mirrors their approach for QED guidance.
@@ B.2.2 Novelty Constraints
For this application, we similarly follow the FUDGE implementation and configuration in [schiff2024simple] for QED guidance.
@@
We use classifier based guidance and classifier free guidance for QED as implemented in [schiff2024simple].
B.3 Instruction Following
@@
To quantify the fluency and coherence of generated continuations we follow the increasingly adopted LLM-as-a-Judge evaluation. This method has been shown to result in similar evaluation as human judgment [zheng2023judgingllmasajudgemtbenchchatbot]. Specifically, each post-prefix sentence is scored by the instruction-tuned Gemini-2.0-Flash model and issued the following prompt:
@@
While in Section 5.1, we directly utilize the pretrained MDLM checkpoint provided by [sahoo2024simple], for the experiments in Section 5.3, we pretrain and subsequently instruction-tune a slightly larger MDLM model, adopting the Llama 1 tokenizer. Empirical observations indicated that employing a marginally larger model and vocabulary size was essential for optimal performance on instruction-guided tasks. Consequently, we adopt this enhanced model for all subsequent evaluations in this context, including both the base MDLM experiments and the proposed CDD framework. Our instruction tuning is conducted using the ShareGPT dataset, and all pretraining follows the design choices and hyperparameters provided by [sahoo2024simple].
B.3.1 Counting Constraints
@@ B.3.2 Lexical Constraints
We finetune models on a dataset of examples constructed from evaluation of GPT-4 on the COLLIE dataset provided by yao2023collie. As GPT-4 performs well on these examples, we apply knowledge distillation by compiling these generations into an instruction tuning dataset. The dataset can be directly constructed from the logs included in the code base from yao2023collie
@@
Appendix C Hyper-parameter Ablation Study
@@
We include a constraint‑start ablation, where we examine the runtime/perplexity tradeoff by turning the projection operator on only after specific amounts of reverse‑diffusion steps. Early enforcements allows the projection to satisfy the constraints at very early iterations at the cost of increased runtime. Notably, across all setting, constraint satisfaction is met for all samples.
ALM hyper‑parameters Perplexity Wall‑clock time (s) LLM‑Judge
@@
Table 3: Perplexity and wall‑clock time statistics, together with LLM‑Judge coherence scores, across all ALM hyper‑parameter settings.
@@
$\text{GPT2}_{\text{PPLM}}$ 124M 878.47 857.38 869.69 887.41 85.21
$\text{GPT2}_{\text{FUDGE}_{\lambda=2}}$ 124M 1387.28 1386.47 1387.00 1387.73 134.56
$\text{GPT2}_{\text{FUDGE}_{\lambda=9}}$ 124M 1481.94 1479.71 1480.27 1484.47 143.74
@@
Appendix D Runtime Analysis
To benchmark the computational overhead of CDD as compared to the baselines, we compare the runtime of our method and the other baselines on the toxicity mitigation task. To normalize for the overhead of base models (MDLM and GPT2), we make two adjustment: we reduce the number of diffusion steps from 1000 to 100 and adjust the generation sequence length to 1024; prior work on discrete diffusion has indicated that these hyperparameters can be used to approximately match the inference time of autoregressive models [loudiscrete]. Empirically, we find that, while the speed is not identical, these adjustments result in a much more similar tokens-per-second rate for the base models. Practically, MDLM could be accelerated even further by reducing the number of diffusion steps further, but we abstain from this so as to maintain consistency with the baselines. For every system we report the mean generation time, the 25th/50th/75th percentiles, and the slowdown relative to the respective baseline.
Appendix E Entropy Analysis
For the toxicity application, we additionally evaluate CDD and its unconstrained base model using entropy. As described in [zheng2024masked], entropy measures the diversity of the generated sequence. For a generated sequence of length $L$ with  $K$ distinct tokens, let $L_{k}$ denote the count of token $k$, and define the token probabilities $p_{k}=L_{k}/L$. The entropy of this distribution is $\sum_{k=1}^{K}p_{k}\log p_{k}$. Lower entropy indicates that probability mass is concentrated on a few tokens, reflecting a lower diversity, whereas a higher entropy indicates a more uniform token usage and greater generative diversity.
As observed in Table 5, across all thresholds, CDD results in a negligible entropy decrease of 0.59–1.0% compared to the base model. As the decrease is small and within narrow confidence intervals, this validates that CDD preserves generative diversity.
Model Size Mean Entropy 95% CI % Decrease
MDLM 110M 5.577 0.0178
[…]
Table 5: CDD mean entropy, 95% confidence intervals, and percent decrease from base model at different thresholds for the toxicity application.
@@ Appendix F Missing Proofs
Since $\mathbf{C}$ is $\beta\text{-prox}$ regular, by definition the following hold:
@@
since projections onto closed prox-regular sets are $L\text{-Lipshitz}$.
@@
Since $\mathbf{C}$ in $\beta\text{-prox}$ regular, for any point $u$ near $\mathbf{C}$ and $v\in\mathbf{C}$: